what is the equation in point-slope form of the line perpendicular to 3x+6y=5 and through the point (1,3).

Answer by stanbon(75887)

what is the equation in point-slope form of the line perpendicular to 3x+6y=5 and through the point (1,3).
----
Solve for "y":
y = (-1/2)x + (5/6)
slope is -1/2
----
Slope of all perpendicular must be 2
----
Equation Form:
y = mx + b
Using (1,3) and m = 2 solve for "b:
3 = 2*1 + b
b = 1
---------------
Equation:
y = 2x + 1
==================
Cheers,
Stan H.
==================